OpenStudy (anonymous):

total tickets 925 total cash $2495 adult tickets $5 child $2 how many of each were sold

OpenStudy (anonymous):

5x+2y=2495 x+y=925

OpenStudy (anonymous):

substitute x=925-y 5(925-y)+2y=2495

OpenStudy (anonymous):

4625-5y+2y=2495 -3y=-2130 y=710

OpenStudy (anonymous):

Now plug that into 1 of the equations and solve for x x+y=925 x+710=925 x=215

OpenStudy (anonymous):

215 adult tickets, 710 kid tickets

OpenStudy (anonymous):

thanks that was a difficult one!

OpenStudy (anonymous):

welcome! medal?

OpenStudy (anonymous):

thanks that was a difficult one!